Find f(t).
5x-1 Find fit). L{ x2(x+1) f(x)=? Use Laplace.

f(t)= -3t^2 e^-t -7te^-t -t+8 -8e^-t...?
Answer #1

solution: O B £ + + ایم اے + 3 f 2 n² UH Using partial fraction, L-S 51-1 n² (4+)3 Let F (u) = 54-1 n² (4+1)3 51-1 D x² (nt (u, Atc=0 Ple= -81 (48) 8 from equation 6 21+0=-23 -16 to =-23 110=-7 from equation , we get C+D+E=-21 7-8-77E = -21 E = -6 .